A hollow copper (ρCu = 8.960 ✕ 103 kg/m3) spherical shell of mass m = 0.9860 kg floats in water with its entire volume below the surface. (Assume the temperature of the water is 4°C. Enter your answers to at least four significant figures.)
4 5 points KatzPSEf1 15 P025.MI. My Notes Ask Your Teacher A hollow copper (Cu-8.960 x 103 kg/m3) spherical shell of mass m = 0.9860 kg floats in water with its entire volume below the surface. (Assume the temperature of the water is 4°C. Enter your answers to at least four significant figures.) (a) What is the radius of the sphere? (b) What is the thickness of the shell wall? Need Help? Lenaar. Latter.」
Answer #1


(a)


In equilibrium net force = 0

Fb - mg = 0

Fb = m*g


Fb = buoyancy force = rho_water*V*g

V = volume of the sphere = (4/3)*pi*R^3


rho_water*(4/3)*pi*R^3*g = m*g

1000*(4/3)*pi*R^3*9.8 = 0.9860*9.8


radius r = 61.74 mm

=========================

part B


mass = density of copper * volume of copper


volume of copper v = (4/3)*pi*(R^3-r^3)

0.9860 = 8.96*10^3*(4/3)*pi*((61.74*10^-3)^3-r^3)


inner radius of the shell r = 59.35 mm


thickness of the shell = R - r = 2.390 mm <<<----ANSWER